# Provably Correct Automata Embeddings for Optimal Automata-Conditioned Reinforcement Learning

### v1

@@ 3.1 PAC-Learnability of DFA-Conditioned Reinforcement Learning Problem
s.t., with probability at least $1-p$, the total number of time steps during which the policy $\pi$ (current policy being trained) executed by $\mathbb{A}$ is more than $\epsilon$ suboptimal is at most $N$, i.e., we have
@@
If a learning algorithm $\mathbb{A}$ is PAC-MDP as defined in 7, then for any DFA-conditioned MDP $\mathcal{M}\mid\mathcal{M}_{D_{\Sigma}}$, $\epsilon>0$, and $p\in(0,1)$, there exists a polynomial function
@@
s.t. the total number of $\epsilon$ suboptimal steps taken by $\mathbb{A}$ is at most $N^{\prime}$ with probability at least $1-p$.
The proof is given in the appendix. 1 proves that the DFA-conditioned RL problem is PAC-learnable, assuming the given MDP is solvable. However, in practice, one cannot input a DFA to a policy directly. Instead, one uses an encoder (possibly pretrained) mapping DFAs to embeddings. In such cases, the optimality of the learned DFA-conditioned policy depends on the encoder.
@@ 4 Learning Provably Correct Automata Embeddings
In the previous section, we introduced the idealized, theoretical formulation of the DFA-conditioned RL problem and proved that it is PAC-learnable. However, a policy implemented by a feed-forward neural network, as is usually the case, cannot condition on a DFA directly, but rather an encoding of the DFA is required, such as a vector representation. In such cases, the question then is whether the policy conditioning on DFA encodings is optimal w.r.t. the theoretical formulation of the problem. This is the problem we tackle in the following, but before doing so, we formally state the problem.
@@
Given a DFA space $D_{\Sigma,n}$, learn an encoder $\phi:D_{\Sigma,n}\to\mathcal{Z}$ s.t. in any underlying MDP $\mathcal{M}$, training a policy $\pi_{\mathcal{Z}}:S\times\mathcal{Z}\to\Delta(A)$ using the latent representations of DFAs in $D_{\Sigma,n}$ converges to the same composite function as a DFA-conditioned policy $\pi:S\times D_{\Sigma,n}\to\Delta(A)$ , formally
.$\forall\mathcal{M}.\quad\arg\max_{\pi}J_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}(\pi)=\arg\max_{\pi_{\mathcal{Z}}\circ\phi}J_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}(\pi_{\mathcal{Z}}\circ\phi),$
@@
$\phi:D_{\Sigma,n}\to\mathcal{Z}$ has enough capacity to represent DFAs in its domain. That is, the learnable encoder $\phi$ has a parametrization that can map distinct DFAs in $D_{\Sigma}$ to unique embeddings.
@@ 4.1 Bisimulation Relations and Metrics over MDP states
A bisimulation metric can be viewed as a relaxation of a notion of bisimulation relation over MDP states. We start by defining the latter and then continue with the former.
@@
$\forall\mathcal{A},\mathcal{A}^{\prime}\in\mathcal{D}_{\Sigma,n}.\quad\mathcal{A}\sim\mathcal{A}^{\prime}\iff\mathcal{A}\sim_{\mathcal{M}_{D_{\Sigma,n}}}\mathcal{A}^{\prime}.$
The proof is given in the appendix. Essentially, 1 shows that 3 and 8 are equivalent for DFAs in a DFA space, which will later help us with 1. We will continue with the formal definition of a bisimulation metric. But before doing so, informally, a function $d:X\times X\to\mathbb{R}_{\geq 0}$ is called a pseudometric on a set $X$ if it satisfies non-negativity, symmetry, and the triangle inequality but may allow $d(x,y)=0$ for $x\neq y$, see the appendix for a formal definition.
@@
=$\sim_{\mathcal{M}}=\{(s,t)\in S\times S\mid d(s,t)=0\}.$
Recall that a given DFA space $D_{\Sigma,n}$ induces a deterministic MDP $\mathcal{M}_{D_{\Sigma,n}}$, where each state of this MDP is a DFA. We want our learned encoder to uniquely distinguish between different behaviors, but we do not care whether we can distinguish between different representations of the same task. Therefore, we can use a bisimulation metric to measure how bisimilar two DFAs are and utilize this idea to learn a provably correct embedding space by ensuring that if two DFAs are not bisimilar, then they must have different embeddings. To this end, we first present the following result stating that bisimulation metrics over deterministic MDPs can be computed as a fixed point solution.
@@ 4.2 Learning Automata Embeddings by Computing Bisimulation Metrics
$a_{t}\sim(\pi\circ\phi)(\mathcal{A}_{t},\mathcal{A}_{t}^{\prime})=\pi(\phi(\mathcal{A}_{t}),\phi(\mathcal{A}_{t}^{\prime})),\quad\mathcal{A}_{t+1}=T_{D_{\Sigma,n}}(\mathcal{A}_{t},a_{t}),\quad\mathcal{A}_{t+1}^{\prime}=T_{D_{\Sigma,n}}(\mathcal{A}_{t}^{\prime},a_{t}),$
@@
$J_{D_{\Sigma}}(\pi\circ\phi)=J_{\text{clip}}(\pi\circ\phi)+J_{\text{val}}(\phi),$ (2)
@@
where $\bar{V}_{t+1}$ and $\bar{d}(\mathcal{A}_{t+1},\mathcal{A}_{t+1}^{\prime})$ denotes calls with stop gradients, i.e., no gradient flow to $\phi$. $J_{\text{val}}(\phi)$ implements the objective for the pseudometric given in 2, penalizing for diverging from the one-step lookahead target. The combined objective of $\phi$ is then to learn latent representations that form a bisimulation metric under normalized $\ell_{2}$-norm while also providing representations for $\pi$. Next, we show that an encoder solving Equation 2 maps two DFAs to the same embedding if and only if they are bisimilar and therefore proving that such encoders can distinguish distinct tasks.
@@
$\forall\mathcal{A},\mathcal{A}^{\prime}\in D_{\Sigma,n}.\quad\mathcal{A}\sim\mathcal{A}^{\prime}\iff\phi^{*}(\mathcal{A})=\phi^{*}(\mathcal{A}^{\prime}).$
The proof is given in the appendix. Observe that if our trained encoder can distinguish between DFAs that are not bisimilar, then it solves 1, as bisimilar DFAs are different representations for the same task–no need to distinguish them. Next, we formally state this result solving 1.
@@ 5 Experiments
We implemented the technique given in Section 4 using a GATv2 model as our DFA encoder and Reach-Avoid Derived (RAD) DFAs with at most $10$ states, which are plan DFAs, both presented in [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]. One difference in our GATv2 model is that given a DFA with $n$ states, we do $n$ message-passing steps, instead of doing it for a fixed number as in [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]. To break the symmetry, caused by taking the absolute value of the reward difference, we trained the policy using the reward difference without the absolute value. LABEL:fig:bisim-curves shows that our training technique finds the fixed point, where the objectives from Section 4 are given as losses. We then tested the accuracy as well as the generalization capabilities of these DFA embeddings. To do so, we generated RAD, Reach (R), and Reah-Avoid (RA) DFAs. During training the number of states of a RAD DFA was sampled from a truncated geometric distribution (with $10$ as the upper bound) whereas during testing we sampled it using a bounded uniform distribution. We also generated out-of-distribution (OOD) DFAs with the number of states sampled uniformly between $11$ and $20$. We computed bisimulation metrics, i.e., the normalized $\ell_{2}$-norms, between the embeddings of these DFAs. LABEL:fig:heatmap gives these results in the form of a heatmap, demonstrating the correctness of the learned DFA embeddings–00 on the diagonal. We further checked whether any of these sampled DFAs (both in-distribution and OOD ones) are mapped to the same embedding (up to a $10^{-8}$ accuracy) or not, and we confirm that the encoder has a $100\%$ success rate in these samples. Finally, LABEL:fig:policy compares our new pretraining technique from Section 4 with our previous pretraining procedure based on solving DFAs ([Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]) which does not guarantee correctness of the learned embeddings. LABEL:fig:policy shows that the correctness of DFA embeddings improves learning.
@@ Appendix A Goal-Conditioned Reinforcement Learning
For reference, here, we define the standard goal-conditioned reinforcement learning (GCRL) problem. We start by defining the environment model for conditioning on goals, where goals are usually defined as sets of states or continuous regions, as done by [Schaul et al.(2015)Schaul, Horgan, Gregor, and Silver], [Liu et al.(2022)Liu, Zhu, and Zhang].
@@
Given a goal-conditioned MDP, the standard goal-conditioned reinforcement learning (GCRL) problem is to find a policy accomplishing a given goal, introduced by [Schaul et al.(2015)Schaul, Horgan, Gregor, and Silver].
@@ Appendix C Proofs of Theorems and Lemmas
A DFA-conditioned MDP $\mathcal{M}\mid\mathcal{M}_{D_{\Sigma}}$ is defined over the state space $S\times D_{\Sigma,n}$, where $D_{\Sigma,n}$ is a DFA space. Since $D_{\Sigma,n}$ is finite (as all DFAs in $D_{\Sigma,n}$ has a finite alphabet $\Sigma$ and at most $n$ states), the product state space has size $|S|\cdot|D_{\Sigma,n}|$ and is an MDP. Therefore, any PAC-MDP algorithm that works for MDPs with state space size $|S|$ will also work on the product MDP with state space size $|S|\cdot|D_{\Sigma,n}|$, with sample complexity increasing by at most a factor polynomial in $|D_{\Sigma,n}|$.

### v2

@@ 3.1 PAC-Learnability of the DFA-Conditioned Reinforcement Learning Problem
s.t., with probability at least $1-p$, the total number of time steps during which the policy $\pi$ (current policy being trained) executed by $\mathbb{A}$ is more than $\epsilon$-suboptimal is at most $N$, i.e., we have
@@
If a learning algorithm $\mathbb{A}$ is PAC-MDP as defined in 7, then for any DFA-conditioned MDP $\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}$, $\epsilon>0$, and $p\in(0,1)$, there exists a polynomial function
@@
s.t. the total number of $\epsilon$-suboptimal steps taken by $\mathbb{A}$ is at most $N^{\prime}$ with probability at least $1-p$.
The proof is in the appendix. 1 proves that the DFA-conditioned RL problem is PAC-learnable, assuming the underlying MDP is solvable. However, in practice, one cannot input a DFA to a policy directly. Instead, one uses an encoder (possibly pretrained) mapping DFAs to embeddings. In such cases, the optimality of the learned DFA-conditioned policy depends on the encoder.
@@ 4 Learning Provably Correct Automata Embeddings
In the previous section, we introduced the idealized, theoretical formulation of the DFA-conditioned RL problem and proved that it is PAC-learnable. However, a policy implemented by a feed-forward neural network, as is usually the case, cannot condition on a DFA directly, but rather an encoding of the DFA is required, such as a vector representation. Then, the question is whether the policy conditioning on DFA encodings is optimal w.r.t. the theoretical formulation of the problem. This is the problem we tackle in the following, but before doing so, we formally state the problem.
@@
Given a DFA space $D_{\Sigma,n}$, learn an encoder $\phi:D_{\Sigma,n}\to\mathcal{Z}$ s.t. for any MDP $\mathcal{M}$ and labeling function $L$, solving $\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}$ with a policy $\pi_{\mathcal{Z}}:S\times\mathcal{Z}\to\Delta(A)$ conditioning on DFA embeddings is equivalent to solving it with a DFA-conditioned policy $\pi:S\times D_{\Sigma,n}\to\Delta(A)$, i.e.,
$\forall\mathcal{M},\forall L,\quad\arg\max_{\pi}J_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(\pi)=\arg\max_{\pi_{\mathcal{Z}}\circ\phi}J_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(\pi_{\mathcal{Z}}\circ\phi),$
@@
$\phi:D_{\Sigma,n}\to\mathcal{Z}$ has enough capacity to represent DFAs in its domain. That is, the learnable encoder $\phi$ has a parametrization that can map distinct DFAs in $D_{\Sigma,n}$ to unique embeddings.
@@ 4.1 Bisimulation Relations and Metrics over MDP states
A bisimulation metric can be viewed as a relaxation of the notion of a bisimulation relation over MDP states. We start by defining the latter and then continue with the former.
@@
$\forall\mathcal{A},\mathcal{A}^{\prime}\in\mathcal{D}_{\Sigma,n},\quad\mathcal{A}\sim\mathcal{A}^{\prime}\iff\mathcal{A}\sim_{\mathcal{M}_{D_{\Sigma,n}}}\mathcal{A}^{\prime}.$
The proof is given in the appendix. Essentially, 1 shows that 3 and 8 are equivalent for DFAs in a DFA space, which will later help us with 1. We will continue with the formal definition of a bisimulation metric. But before doing so, informally, a function $d:X\times X\to\mathbb{R}_{\geq 0}$ is called a pseudometric on a set $X$ if it satisfies non-negativity, symmetry, and the triangle inequality but may allow $d(x,y)=0$ for $x\neq y$, see Appendix B for a formal definition.
@@
=$\sim_{\mathcal{M}}=\{(s,t)\in S\times S\mid_{L}d(s,t)=0\}.$
Recall that a given DFA space $D_{\Sigma,n}$ induces a deterministic MDP $\mathcal{M}_{D_{\Sigma,n}}$, where each state of this MDP is a DFA. We want our learned encoder to uniquely distinguish between different behaviors, but we do not care whether we can distinguish between different representations of the same task. Therefore, we can use a bisimulation metric to measure how bisimilar two DFAs are and utilize this idea to learn a provably correct embedding space by ensuring that if two DFAs are not bisimilar, then they must have different embeddings. To this end, we first present the following result stating that bisimulation metrics over deterministic MDPs can be computed as fixed-point solutions.
@@ 4.2 Learning Automata Embeddings by Computing Bisimulation Metrics
$a_{t}\sim(\pi\circ\phi)(\mathcal{A}_{t},\mathcal{A}_{t}^{\prime}),\quad\mathcal{A}_{t+1}=T_{D_{\Sigma,n}}(\mathcal{A}_{t},a_{t}),\quad\mathcal{A}_{t+1}^{\prime}=T_{D_{\Sigma,n}}(\mathcal{A}_{t}^{\prime},a_{t}),$
@@
$J_{D_{\Sigma,n}}(\pi\circ\phi)=J_{\text{clip}}(\pi\circ\phi)+J_{\text{val}}(\phi),$ (2)
@@
where $\bar{V}_{t+1}$ and $\bar{d}(\mathcal{A}_{t+1},\mathcal{A}_{t+1}^{\prime})$ denotes calls with stop gradients, i.e., no gradient flow to $\phi$. $J_{\text{val}}(\phi)$ implements the objective for the pseudometric given in 2, penalizing for diverging from the one-step lookahead target. The combined objective of $\phi$ is then to learn latent representations that form a bisimulation metric under normalized $\ell_{2}$-norm while also providing representations for $\pi$. Next, we show that an encoder maximizing Equation 2 maps two DFAs to the same embedding if and only if they are bisimilar, therefore proving that such encoders can distinguish distinct tasks.
@@
$\forall\mathcal{A},\mathcal{A}^{\prime}\in D_{\Sigma,n},\quad\mathcal{A}\sim\mathcal{A}^{\prime}\iff\phi^{*}(\mathcal{A})=\phi^{*}(\mathcal{A}^{\prime}).$
The proof is given in the appendix. Observe that if our trained encoder can distinguish between DFAs that are not bisimilar, then it solves 1, as bisimilar DFAs are different representations for the same task–no need to distinguish them. Next, we formally state this result, solving 1.
@@ 5 Experiments
We implemented the technique given in Section 4 using a GATv2 model as our DFA encoder and Reach-Avoid Derived (RAD) DFAs with at most $10$ states, which are plan DFAs, both presented in [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]. One difference in our GATv2 model is that given a DFA with $n$ states, we do $n$ message-passing steps, instead of doing it for a fixed number as in [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]. To break the symmetry, caused by taking the absolute value of the reward difference, we trained the policy using the reward difference without the absolute value. LABEL:fig:bisim-curves shows that our training technique finds the fixed point, where the objectives from Section 4 are given as losses. We then tested the accuracy as well as the generalization capabilities of these DFA embeddings. To do so, we generated RAD, Reach (R), and Reah-Avoid (RA) DFAs. During training the number of states of a RAD DFA was sampled from a truncated geometric distribution (with $10$ as the upper bound) whereas during testing we sampled it using a bounded uniform distribution. We also generated out-of-distribution (OOD) DFAs with the number of states sampled uniformly between $11$ and $20$. We computed bisimulation metrics, i.e., the normalized $\ell_{2}$-norms, between the embeddings of these DFAs. LABEL:fig:heatmap gives these results in the form of a heatmap, demonstrating the correctness of the learned DFA embeddings–00 on the diagonal. We further checked whether any of these sampled DFAs (both in-distribution and OOD ones) are mapped to the same embedding (up to a $10^{-8}$ accuracy) or not, and we confirm that the encoder has a $100\%$ success rate in these samples. LABEL:fig:policy compares our new pretraining technique from Section 4 with our previous pretraining procedure based on solving DFAs ([Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]), which does not guarantee correctness, showing that the correctness of DFA embeddings improves downstream policy learning. All results are over 5 seeds.
@@ Appendix A Goal-Conditioned Reinforcement Learning
Here, for reference, we present the standard Goal-Conditioned Reinforcement Learning (GCRL) problem. We start by defining the environment model for conditioning on goals, usually given as sets of states or continuous regions, as done by [Schaul et al.(2015)Schaul, Horgan, Gregor, and Silver, Liu et al.(2022)Liu, Zhu, and Zhang].
@@
Given a goal-conditioned MDP, the standard GCRL problem is to find a policy that achieves a given goal, which was first introduced by [Schaul et al.(2015)Schaul, Horgan, Gregor, and Silver].
@@ Appendix C Proofs of Theorems and Lemmas
A DFA-conditioned MDP $\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}$ is defined over the state space $S\times D_{\Sigma,n}$, where $D_{\Sigma,n}$ is a DFA space. Since $D_{\Sigma,n}$ is finite (as all DFAs in $D_{\Sigma,n}$ has a finite alphabet $\Sigma$ and at most $n$ states), the product state space has size $|S|\cdot|D_{\Sigma,n}|$ and is an MDP. Thus, any PAC-MDP algorithm that works for MDPs with state space size $|S|$ will also work on the product MDP with state space size $|S|\cdot|D_{\Sigma,n}|$, with sample complexity increasing by at most a factor polynomial in $|D_{\Sigma,n}|$.